# Revisiting Few-Shot Object Detection with Vision-Language Models

### v1

@@ 4 Experiments
  • nuImages [1] annotates 18 classes, which are divided into groups with many, medium, and few examples [29, 26] and report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. Unlike LVIS, nuImages is fully annotated, so we construct our Foundational FSOD benchmark using the procedure described in subsection $3.$ Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 10 random data splits and report the average.
@@ 4.2 Foundational FSOD with LVIS
Table 3: nuImages Foundational FSOD Performance. We repurpose nuImages for FSOD following the dataset creation process established by [38]. We group categories by frequency into cohorts with many, medium and few examples per class (according to the fully-annotated dataset) [29, 26]. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples for each of the 18 nuImages classes. We observe that existing methods like TFA perform poorly on nuImages, achieving <3<3AP. However, by simply pre-training TFA on more data and leveraging language cues via CLIP classifier, we dramatically improve performance (from 2.0 AP to 14.8 AP, 5-shots). Interestingly, accuracy across cardinalities decreases despite all classes being trained with $K$ examples. This suggests that despite pre-training on web-scale datasets, VLMs still struggle to detect rare categories like strollers, pushable-pullable, and debris, highlighting the challenge of working with nuImages.
@@ 4.3 Foundational FSOD with nuImages
To contextualize our results, we evaluate TFA [38] on the nuImages Foundational FSOD benchmark. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than 33 AP (cf. Table 3). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
@@ Appendix A Implementation Details
LVIS v0.5 Experiment Details. We select Detic with a Resnet-50 backbone for fair comparison with prior work. We pre-train Detic on LVIS-base for $90k$ iterations with a batch size of $32$ using an AdamW optimizer and a learning rate of 2e32e-3. All images are resized to $640\times 640$ and we also enable Repeat Factor Sampling [11].
Following [38], we sample up to $10$ shots for each class in LVIS (since all classes may not have 10 examples). We use a batch size of $32$, learning rate of 2.5e52.5e-5 for $46k$ iterations. We do not use Repeat Factor Sampling for fine-tuning. We sample $50$ categories for each training image, i.e $|S|=50$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
nuImages Experiment Details. We select Detic with a Swin-T backbone pre-trained on LVIS+COCO and ImageNet-21k data. We use an image size of $1600\times 900$, batch size of 88 and an AdamW optimizer with learning rate of 3.75e63.75e-6. We fine-tune this model for $8000$ iterations on nuImages. We sample 66 categories for each training image, i.e $|S|=6$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
@@ Appendix B Analysis of Iconic Few-Shot Images
We construct our best split by picking examples corresponding to the best class-wise performance, based on the evaluation of each split on a held-out validation set. For instance, out of $10$ random splits for the 55-shot task, one may pick car examples from split 11, bicycle from split 44 and debris from split 88. In Table 6, we observe that the best-split performance is always better than its random counterpart. As expected, the choice of examples in 55-shot case is more important than the $30$-shot case ($1.63$ AP difference for 55-shot vs $0.08$ AP for $30$-shots). We visualize the difference in the splits for the stroller class in nuImages (cf. Figure 5). Unsurprisingly, iconic examples are large and unoccluded.
@@
Table 6: Random Split vs ‘Best” Split. We construct the “best” split by selecting per-class few-shot examples that lead to the highest performance on a held out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. 55-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
@@
Naive fine-tuning improves over zero shot results by $\sim 2.5$ AP ($16.83$ vs $14.26$ AP) in the $K=30$ shot setting. In addition, we see diminishing benefits for using FedLoss variants (cf. Table 8). Nevertheless, pseudo-negatives improve over naive fine-tuning by 11 AP ($16.83$ vs $17.87$ AP). Unsurprisingly, performance for classes with many examples increases the most ($19.40$ AP) when training with exhaustive annotations. Importantly, training with exhaustive annotations uses orders of magnitude more examples (1300 examples of car) than with few-shot annotations (30 examples of car).
Table 7: Analysis of 55-shot Performance on nuImages. Fine-tuning with pseudo-negatives improves over naive fine-tuning. Importantly, our pseudo-negative approach nearly matches the upper bound with true negatives ($15.94$ AP vs $16.15$ AP). This suggests that our proposed approach is particularly effective in data-constrained settings.
@@ Appendix D RegionCLIP Experiments
We report RegionCLIP’s zero-shot and fine-tuning performance on nuImages averaged over 33 random splits in Table 9. Detic zero-shot outperforms RegionCLIP zero-shot by $\sim 12$ AP ($14.26$ vs $2.34$). While fine-tuning RegionCLIP improves overall performance, Detic achieves higher accuracy for $K=\{5,10,30\}$ shots. This highlights the importance of supervision type (e.g. box-supervised data) and data scale used for pre-training.
@@
Lastly, we evaluate RegionCLIP’s performance with Detic-RPN. Notably, we observe that the performance improves over RegionCLIP w/ LVIS-RPN demonstrating that reducing the number of false positive proposals yields better performance. Furthermore, we filter out low confidence Detic proposals , i.e <0.5<0.5 objectness score (w/ Detic-RPN, 0.5) and find that this doubles RegionCLIP’s zero-shot performance to $7.64$ AP.

### v2

@@ 4 Experiments
  • nuImages [1] annotates 18 classes, which are divided into groups with many, medium, and few examples [29, 26] and report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. Unlike LVIS, nuImages is fully annotated, so we construct our Foundational FSOD benchmark using the procedure described in subsection 3.3. Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 10 random data splits and report the average.
@@ 4.2 Foundational FSOD with LVIS
Table 3: nuImages Foundational FSOD Performance. We repurpose nuImages for FSOD following the dataset creation process established by [38]. We group categories by frequency into cohorts with many, medium and few examples per class (according to the fully-annotated dataset) [29, 26]. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples for each of the 18 nuImages classes. We observe that existing methods like TFA perform poorly on nuImages, achieving $<3$AP. However, by simply pre-training TFA on more data and leveraging language cues via CLIP classifier, we dramatically improve performance (from 2.0 AP to 14.8 AP, 5-shots). Interestingly, accuracy across cardinalities decreases despite all classes being trained with $K$ examples. This suggests that despite pre-training on web-scale datasets, VLMs still struggle to detect rare categories like strollers, pushable-pullable, and debris, highlighting the challenge of working with nuImages.
@@ 4.3 Foundational FSOD with nuImages
To contextualize our results, we evaluate TFA [38] on the nuImages Foundational FSOD benchmark. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than $3$ AP (cf. Table 3). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
@@ Appendix A Implementation Details
LVIS v0.5 Experiment Details. We select Detic with a Resnet-50 backbone for fair comparison with prior work. We pre-train Detic on LVIS-base for $90k$ iterations with a batch size of $32$ using an AdamW optimizer and a learning rate of $32e-$ 3. All images are resized to $640\times 640$ and we also enable Repeat Factor Sampling [11].
Following [38], we sample up to $10$ shots for each class in LVIS (since all classes may not have 10 examples). We use a batch size of $32$, learning rate of $52.5e-$ 5 for $46k$ iterations. We do not use Repeat Factor Sampling for fine-tuning. We sample $50$ categories for each training image, i.e $|S|=50$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
nuImages Experiment Details. We select Detic with a Swin-T backbone pre-trained on LVIS+COCO and ImageNet-21k data. We use an image size of $1600\times 900$, batch size of $8$ and an AdamW optimizer with learning rate of $63.75e-$ 6. We fine-tune this model for $8000$ iterations on nuImages. We sample $6$ categories for each training image, i.e $|S|=6$ for the FedLoss and InvFedLoss experiments. We derive negatives from pseudolabels with atleast $20\%$ confidence for the Psuedo-Negative experiment.
@@ Appendix B Analysis of Iconic Few-Shot Images
We construct our best split by picking examples corresponding to the best class-wise performance, based on the evaluation of each split on a held-out validation set. For instance, out of $10$ random splits for the $5$-shot task, one may pick car examples from split $1$, bicycle from split $4$ and debris from split $8$. In Table 6, we observe that the best-split performance is always better than its random counterpart. As expected, the choice of examples in $5$-shot case is more important than the $30$-shot case ($1.63$ AP difference for $5$-shot vs $0.08$ AP for $30$-shots). We visualize the difference in the splits for the stroller class in nuImages (cf. Figure 5). Unsurprisingly, iconic examples are large and unoccluded.
@@
Table 6: Random Split vs ‘Best” Split. We construct the “best” split by selecting per-class few-shot examples that lead to the highest performance on a held out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. $5$-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
@@
Naive fine-tuning improves over zero shot results by $\sim 2.5$ AP ($16.83$ vs $14.26$ AP) in the $K=30$ shot setting. In addition, we see diminishing benefits for using FedLoss variants (cf. Table 8). Nevertheless, pseudo-negatives improve over naive fine-tuning by $1$ AP ($16.83$ vs $17.87$ AP). Unsurprisingly, performance for classes with many examples increases the most ($19.40$ AP) when training with exhaustive annotations. Importantly, training with exhaustive annotations uses orders of magnitude more examples (1300 examples of car) than with few-shot annotations (30 examples of car).
Table 7: Analysis of $5$-shot Performance on nuImages. Fine-tuning with pseudo-negatives improves over naive fine-tuning. Importantly, our pseudo-negative approach nearly matches the upper bound with true negatives ($15.94$ AP vs $16.15$ AP). This suggests that our proposed approach is particularly effective in data-constrained settings.
@@ Appendix D RegionCLIP Experiments
We report RegionCLIP’s zero-shot and fine-tuning performance on nuImages averaged over $3$ random splits in Table 9. Detic zero-shot outperforms RegionCLIP zero-shot by $\sim 12$ AP ($14.26$ vs $2.34$). While fine-tuning RegionCLIP improves overall performance, Detic achieves higher accuracy for $K=\{5,10,30\}$ shots. This highlights the importance of supervision type (e.g. box-supervised data) and data scale used for pre-training.
@@
Lastly, we evaluate RegionCLIP’s performance with Detic-RPN. Notably, we observe that the performance improves over RegionCLIP w/ LVIS-RPN demonstrating that reducing the number of false positive proposals yields better performance. Furthermore, we filter out low confidence Detic proposals , i.e $<0.5$ objectness score (w/ Detic-RPN, 0.5) and find that this doubles RegionCLIP’s zero-shot performance to $7.64$ AP.